Reports 1-1 of 1 Clear search Modify search
koki.okutomi - 20:21 Tuesday 21 May 2019 (8928) Print this report
Satellite box replacement for PR2 IM OSEMs


Satellite box replacement

  • I replaced the satellite boxes for PR2 IM OSEMs to modified ones.
    • See also: klog #8912
    • The old ones have been returned to AEL staffs.
    • Serial numbers of the related satellite boxes
        Old New
      IM1 (V1, V2, V3, H1) S1604896 S1604902
      IM2 (H2, H3) S1604898 S1604890
    • At this time, I initially replaced the IM1 satellite box to S1807633 but the ch1 seems not working.
      • I localized the cause of this problem to this satellite box by swapping the cables with one in another channel.
  • I measured the spectra of the IM OSEMs with both the old and new satellite boxes.
    • All the channels of the new satellite boxes look working without any malfunction.
    • The OSEMs with the new satellite box showed smaller noise levels rather at all frequencies (not only higher frequencies).
      • I have no idea why such improvement can happen with just an additional capacitor.
      • The amplification at the DC shows transient behavior after the cable connection.

Update of the wiki information about the PR racks

BO implementation for remote switching of the stepper motor drivers

  • Beside the wiki updating, I checked the current setup of the BO card and its downstreams.
    • I found the BIO-BNC converter is already mounted on the rack, but they need to be cared about the short between the BNC shield and Dsub37 shell by putting jumper sleeves.
    • The cables to connect the BIO-BNC converter with the stepper motor driver are already installed.
    • I added a BO card on the I/O chassis in the PRM rack.

Next steps

  • Check of the BO functionality with short-cared BIO-BNC converters
  • Position adjustment of the horizonrtal OSEMs on the PR2 IM with the traversers
Search Help